40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 10 19:09:48 2009, Jenne, Update, PSL, PMC LO Calibration RFSliderAdjustCalib.pngRFSliderAdjustCalibWithOsc.png
    Reply  Sat Apr 11 00:45:41 2009, Yoichi, Update, PSL, PMC LO Calibration 
       Reply  Sun Apr 12 19:27:20 2009, rana, Update, PSL, PMC LO Calibration 
       Reply  Mon Apr 13 08:59:57 2009, steve, Update, PSL, mixers on order 
Message ID: 1477     Entry time: Mon Apr 13 08:59:57 2009     In reply to: 1473
Author: steve 
Type: Update 
Category: PSL 
Subject: mixers on order 



I then changed the RF Output Adjust slider in increments of 0.5, and measured the peak-to-peak values on the scope. In the table and on the plots, I've taken into account the 12dB attenuation. i.e I actually measured 964mV, so 964mV*10^.6 = 3838mV.

3.8Vpp is about 16dBm.
The mixer for the PMC demodulator is level 23. So 16dBm is insufficient.
What is the level of the new mixer Steve ordered ? 13 ?

I ordered mixers level 13, 17 on Friday and level 23 now.
They should be here Tuesday

NOTE: level 23 power is illegal to use in the 40m lab
They get hot
ELOG V3.1.3-